# Constructing Safety Cases for AI Systems: A Reusable Template Framework

### v1

@@ Appendix B   List of Selected Studies
  • [S68] Christoph Schmittner, Zhendong Ma, and Thomas Gruber. Model-based security assurance cases for open and adaptive cyber- physical systems. In International Conference on Computer Safety, Reliability, and Security, pages 251–262. Springer, 2018.
  • [S69] Barbara Gallina, Irfan Sljivo, and S Ramesh. Model-connected safety cases. In International Conference on Computer Safety, Reliability, and Security, pages 294–306. Springer, 2014.
  • [S70] Peter Bishop and Robin Bloomfield. Moving towards goal-based safety management. In Safety-Critical Systems Symposium, pages 1–15, 1998.
  • [S71] Julian Münch, Manfred Broy, Stefan Kugele, Christian Karcher, and Michael Vierhauser. nn-dependability-kit: Engineering neural networks for safety-critical autonomous driving systems. preprint arXiv:2201.06409, 2022.
  • [S72] Eric Griffor and Stefan Salinger. Pattern-based safety development methods: Overview and comparison. preprint arXiv:2405.00217, 2024.
  • [S73] Ben Bennett, Zoe Porter, Ibrahim Habli, Tim Kelly, and John McDermid. Principled safety assurance arguments. AI and Ethics, 3(1):51–72, 2023.
  • [S74] Mithila Sivakumar, Alvine Boaye Belle, Jinjun Shan, and Kimya Khakzad Shahandashti. Prompting gpt–4 to support automatic safety case generation. In IEEE/ACM 4th International Conference on AI Engineering—Software Engineering for AI (CAIN), pages 298–309. IEEE, 2025.
  • [S75] Adrian Herman, Jeff Gray, Dipanjan Dey, and Robyn Lutz. Querying automotive system models and safety artifacts with mmint and viatra. In International Conference on Computer Safety, Reliability, and Security, pages 183–194. Springer, 2019.
  • [S76] Simon Burton, Ibrahim Habli, and Timothy Lawton. Reliability assessment and safety arguments for machine learning components in system assurance. Safety Science, 157:105932, 2022.
@@
  • [S77] Frances R Ward, Ibrahim Habli, and Tim Kelly. Safe autonomous systems in a changing world: Operationalising dynamic safety cases. Reliability Engineering & System Safety, 234:109099, 2024.
  • [S78] Eirik Nielsen and Tim Kelly. Safety assurance argument context for automated driving. In International Conference on Computer Safety, Reliability, and Security, pages 17–28. Springer, 2018.
  • [S79] Shun Miyoshi, Takuma Tsuchida, Ryuichi Shima, Hironori Washizaki, and Naoyasu Ubayashi. Safety assurance case description method for systems incorporating off‐operational machine learning and safety device. In International Conference on Computer Safety, Reliability, and Security, pages 169–184. Springer, 2023.
  • [S80] James Hurd, Andrew Boyd, James Flynn, Adam Dodd, and Tim Kelly. Safety assurance of machine learning for autonomous systems. Journal of Systems and Software, 212:112010, 2024.
  • [S81] Sebastian von Essen and Jürgen Mottok. Safety assurance of machine learning for chassis control functions. preprint arXiv:2310.15993, 2023.
  • [S82] Simon Burton, Giacomo Maccabiani, Carlo Picardi, Ram Potham, Julian Reich, Richard Heinrich, and Ernest Wozniak. Safety assurance of machine learning for perception functions. SAE International Journal of Connected and Automated Vehicles, 6(13-06-03-0013), 2023.
@@
  • [S83] Richard Palin and Tim Kelly. Safety case construction and reuse using patterns. In International Conference on Computer Safety, Reliability, and Security, pages 137–150. Springer, 2004.
  • [S84] Jose Luis de la Vara, Salome Maro, Giulia Tagliaferri, Mehrdad Sabetzadeh, and Lionel Briand. Safety case maintenance: A systematic literature review. Information and Software Technology, 132:106482, 2021.
@@
  • [S85] Safety case template for inability arguments. Technical report, AISI UK, 2024.
  • [S86] Zoe Porter, Ibrahim Habli, Marten Kaas, and John McDermid. Safety case template for frontier ai: A cyber inability argument. AI and Ethics, 2025.
  • [S87] Ewen Denney and Ganesh Pai. Safety case templates for autonomous systems. NASA Technical Reports, 2022.
  • [S88] Richard Hawkins and Tim Kelly. Safety cases for advanced control software: Safety case patterns. Reliability Engineering & System Safety, 179:47–65, 2018.
  • [S89] David Krueger, Joshua Clymer, Marten Kaas, Zoe Porter, and Ibrahim Habli. Safety cases for frontier ai. preprint arXiv:2407.16717, 2024.
  • [S90] Zoe Porter, Ibrahim Habli, John McDermid, and Marten Kaas. Safety cases: A scalable approach to frontier AI safety. AI and Ethics, 2025.
@@
  • [S91] Joshua Clymer, Nick Gabrieli, David Krueger, and Thomas Larsen. Safety cases: How to justify the safety of advanced AI systems. preprint arXiv:2403.10462, 2024.
  • [S92] David Krueger, Marten Kaas, and Zoe Porter. Safety integrity levels for artificial intelligence. preprint arXiv:2406.19088, 2024.
  • [S93] Anjali Gupta, Ganesh Pai, Atanu Basak, and Simon Burton. Safety-driven design of machine learning for sepsis treatment. preprint arXiv:2407.09842, 2024.
  • [S94] Danny Weyns, Bradley Schmerl, Radu Calinescu, and Ilias Gerostathopoulos. Self-adaptation strategies to maintain security assurance cases. IEEE Software, 39(3):60–70, 2022.
  • [S95] Jonas Krook, Yuvaraj Selvaraj, Wolfgang Ahrendt, and Martin Fabian. Tackling uncertainty in safety assurance for machine learning: Continuous argument engineering with attributed tests. preprint arXiv:2406.05130, 2024.
  • [S96] Joshua Clymer, David Krueger, and Marten Kaas. The big argument for AI safety cases. preprint arXiv:2503.02222, 2025.
  • [S97] Zoe Porter, Ibrahim Habli, and John McDermid. The big argument for AI safety cases. AI and Ethics, 2025.
  • [S98] Julian Reich, Simon Burton, Giacomo Maccabiani, Carlo Picardi, and Richard Reich. The missing link: Developing a safety case for perception components in automated driving. preprint arXiv:2311.02763, 2023.
@@
  • [S99] Franziska Wolny, Silvia Vock, Tomasz Holoyad, and Ralf Adler. The need for systematic approaches in risk assessment of safety- critical AI applications in machinery. European Safety and Reliability Conference (ESREL), 2025.
@@
  • [S100] William Anderson and Paul Dempsey. The open autonomy safety case framework. preprint arXiv:2502.03187, 2025.
  • [S101] Richard Hawkins and Ibrahim Habli. The use of safety cases in certification and regulation. Reliability Engineering & System Safety, 196:106753, 2020.
  • [S102] Markus Borg and Elias Sonnsjö Lönegren. Three sketches of asl-4 safety case components. preprint arXiv:2403.18265, 2024.
@@
  • [S103] Andreas Klöckner, Benedikt Müller, Jochen Kreimeier, and Eric Ernst. Toward a harmonized approach—requirement-based structuring of a safety assurance argumentation for automated vehicles. preprint arXiv:2404.01269, 2024.
  • [S104] Saurabh Basu, Muhammad Rahman, Shreyasee Ghosh, and Sundaram Subramanian. Towards developing safety assurance cases for learning-enabled medical cyber-physical systems. In International Conference on Computer Safety, Reliability, and Security, pages 103–118. Springer, 2024.
  • [S105] Simon Diemert, Erin Cyffka, Torin Viger, and Olivia Foster. Towards evaluations-based safety cases for AI scheming. preprint arXiv:2503.00169, 2025.
  • [S106] Jose Luis de la Vara, Rajwinder K Panesar-Walawege, Mehrdad Sabetzadeh, and Lionel Briand. Towards goal-based software safety certification based on prescriptive standards. Information and Software Technology, 56(8):922–940, 2014.
@@
  • [S107] Yueming Zhao, Weiqing Lu, Jun Lin, Yujie Liu, and Yang Liu. Trusta: Reasoning about assurance cases with formal methods and large language models. preprint arXiv:2506.02491, 2025.
  • [S108] Thomas Chowdhury and Ewen Denney. Uncertainty in assurance case pattern for machine learning. preprint arXiv:2402.07912, 2024.
@@
  • [S109] Tomek Korbak, Jacob Pfau, and Geoffrey Irving. Upstream and downstream AI safety: Both on the same river? preprint arXiv:2504.01557, 2025.
  • [S110] Mithila Sivakumar, Alvine Boaye Belle, Jinjun Shan, and Kimya Khakzad Shahandashti. Using gpt-4 turbo to automatically identify defeaters in assurance cases. preprint arXiv:2503.05909, 2025.
@@
  • [S111] Thor Myklebust, Tor Stålhane, Gunnar D Jenssen, and Irene Wærø. Autonomous cars, trust and safety case for the public. In 2020 Annual Reliability and Maintainability Symposium (RAMS), pages 1–6. IEEE, 2020.
@@
  • [S112] John Bragg and Ibrahim Habli. What is acceptably safe for reinforcement learning? Reliability Engineering & System Safety, 240:109252, 2025.

### v3

@@ Appendix B   List of Selected Studies
  • [S68] L. Nascimento, A. L. de Oliveira, R. Villela, E. F. Silva, R. Wei, R. Hawkins, and T. Kelly, “Model-based security assurance cases for open and adaptive cyber-physical systems,” in International Conference on Advanced Information Networking and Applications. Springer, 2025, pp. 326–340.
  • [S69] A. Retouniotis, Y. Papadopoulos, I. Sorokos, D. Parker, N. Matragkas, and S. Sharvia, “Model-connected safety cases,” in International Symposium on Model-Based Safety and Assessment. Springer, 2017, pp. 50–63.
  • [S70] D. H. Becht, “Moving towards goal-based safety management,” in Proceedings of the Australian System Safety Conference-Volume 133, 2011, pp. 19–26.
  • [S71] C.-H. Cheng, C.-H. Huang, and G. Nührenberg, “nn-dependability-kit: Engineering neural networks for safety-critical autonomous driving systems,” in 2019 IEEE/ACM International Conference on Computer-Aided Design (ICCAD). IEEE, 2019, pp. 1–6.
  • [S72] C. Preschern, N. Kajtazovic, A. Höller, and C. Kreiner, “Pattern-based safety development methods: overview and comparison,” in Proceedings of the 19th European Conference on Pattern Languages of Programs, 2014, pp. 1–20.
  • [S73] N. Annable, M. Lawford, R. F. Paige, and A. Wassyng, “Principled safety assurance arguments,” in International Conference on Computer Safety, Reliability, and Security. Springer, 2025, pp. 18–32.
  • [S74] M. Sivakumar, A. B. Belle, J. Shan, and K. K. Shahandashti, “Prompting gpt–4 to support automatic safety case generation,” Expert Systems with Applications, vol. 255, p. 124653, 2024.
  • [S75] A. Di Sandro, S. Kokaly, R. Salay, and M. Chechik, “Querying automotive system models and safety artifacts with mmint and viatra,” in 2019 ACM/IEEE 22nd international conference on model driven engineering languages and systems companion (MODELS-c). IEEE Computer Society, 2019, pp. 2–11.
  • [S76] Y. Dong, W. Huang, V. Bharti, V. Cox, A. Banks, S. Wang, X. Zhao, S. Schewe, and X. Huang, “Reliability assessment and safety arguments for machine learning components in system assurance,” ACM transactions on embedded computing systems, vol. 22, no. 3, pp. 1–48, 2023.
@@
  • [S77] L. Buysse, I. Habli, D. Vanoost, and D. Pissoort, “Safe autonomous systems in a changing world: Operationalising dynamic safety cases,” Safety Science, vol. 191, p. 106965, 2025.
  • [S78] S. Ballingall, M. Sarvi, and P. Sweatman, “Safety assurance concepts for automated driving systems,” SAE International Journal of Advances and Current Practices in Mobility, vol. 2, no. 2020-01-0727, pp. 1528–1537, 2020.
  • [S79] H. Fujino, N. Kobayashi, and S. Shirasaka, “Safety assurance case description method for systems incorporating off-operational machine learning and safety device,” in INCOSE International Symposium, vol. 29. Wiley Online Library, 2019, pp. 152–164.
  • [S80] C. Paterson, R. Hawkins, C. Picardi, Y. Jia, R. Calinescu, and I. Habli, “Safety assurance of machine learning for autonomous systems,” Reliability Engineering & System Safety, vol. 264, p. 111311, 2025.
  • [S81] S. Burton, I. Kurzidem, A. Schwaiger, P. Schleiss, M. Unterreiner, T. Graeber, and P. Becker, “Safety assurance of machine learning for chassis control functions,” in International Conference on Computer Safety, Reliability, and Security. Springer, 2021, pp. 149–162.
  • [S82] S. Burton, C. Hellert, F. Hüger, M. Mock, and A. Rohatschek, “Safety assurance of machine learning for perception functions,” in Deep Neural Networks and Data for Automated Driving: Robustness, Uncertainty Quantification, and Insights Towards Safety. Springer International Publishing Cham, 2022, pp. 335–358.
@@
  • [S83] T. P. Kelly and J. A. McDermid, “Safety case construction and reuse using patterns,” in Safe Comp 97: The 16th International Conference on Computer Safety, Reliability and Security. Springer, 1997, pp. 55–69.
  • [S84] C. Cârlan, B. Gallina, and L. Soima, “Safety case maintenance: a systematic literature review,” in International Conference on Computer Safety, Reliability, and Security. Springer, 2021, pp. 115–129.
@@
  • [S85] A. UK, “Safety case template for inability arguments,” AISI UK, Tech. Rep., 2024. [Online]. Available: https://www.aisi.gov.uk/blog/safety-case-template-for-inability-arguments
  • [S86] A. Goemans, M. D. Buhl, J. Schuett, T. Korbak, J. Wang, B. Hilton, and G. Irving, “Safety case template for frontier ai: A cyber inability argument,” arXiv preprint arXiv:2411.08088, 2024.
  • [S87] R. Bloomfield, G. Fletcher, H. Khlaaf, L. Hinde, and P. Ryan, “Safety case templates for autonomous systems,” arXiv preprint arXiv:2102.02625, 2021.
  • [S88] R. Alexander, M. Hall-May, T. Kelly, and J. A. McDermid, “Safety cases for advanced control software: Final report,” University of York, York, United Kingdom, Final Report, Jun. 2007, approved for public release; distribution unlimited.
  • [S89] M. D. Buhl, G. Sett, L. Koessler, J. Schuett, and M. Anderljung, “Safety cases for frontier ai,” arXiv preprint arXiv:2410.21572, 2024.
  • [S90] B. Hilton, M. D. Buhl, T. Korbak, and G. Irving, “Safety cases: A scalable approach to frontier AI safety,” arXiv preprint arXiv:2503.04744, 2025.
@@
  • [S91] J. Clymer, N. Gabrieli, D. Krueger, and T. Larsen, “Safety cases: How to justify the safety of advanced AI systems,” preprint arXiv:2403.10462, 2024.
  • [S92] S. Diemert, L. Millet, J. Groves, and J. Joyce, “Safety integrity levels for artificial intelligence,” in International Conference on Computer Safety, Reliability, and Security. Springer, 2023, pp. 397–409.
  • [S93] Y. Jia, T. Lawton, J. Burden, J. McDermid, and I. Habli, “Safety-driven design of machine learning for sepsis treatment,” Journal of Biomedical Informatics, vol. 117, p. 103762, 2021.
  • [S94] S. Jahan, A. Marshall, and R. Gamble, “Self-adaptation strategies to maintain security assurance cases,” in 2018 IEEE 12th International Conference on Self-Adaptive and Self-Organizing Systems (SASO). IEEE, 2018, pp. 180–185.
  • [S95] Y. Matsuno, F. Ishikawa, and S. Tokumoto, “Tackling uncertainty in safety assurance for machine learning: continuous argument engineering with attributed tests,” in International Conference on Computer Safety, Reliability, and Security. Springer, 2019, pp. 398–404.
  • [S96] I. Habli, R. Hawkins, C. Paterson, P. Ryan, Y. Jia, M. Sujan, and J. McDermid, “The big argument for AI safety cases,” arXiv preprint arXiv:2503.11705, 2025.
  • [S97] R. Salay, K. Czarnecki, H. Kuwajima, H. Yasuoka, V. Abdelzad, C. Huang, M. Kahn, V. D. Nguyen, and T. Nakae, “The missing link: Developing a safety case for perception components in automated driving,” SAE International Journal of Advances and Current Practices in Mobility, vol. 5, no. 2022-01-0818, pp. 567–579, 2022.
@@
  • [S98] F. Wolny, S. Vock, T. Holoyad, and R. Adler, “The need for systematic approaches in risk assessment of safety-critical ai-applications in machinery,” in European Safety and Reliability & Society for Risk Analysis Europe Conference ESREL SRA-E, 2025.
@@
  • [S99] M. Wagner and C. Carlan, “The open autonomy safety case framework,” arXiv preprint arXiv:2404.05444, 2024.
  • [S100] N. Leveson, “The use of safety cases in certification and regulation,” Journal of System Safety, 2011, aeronautics and Astronautics/Engineering Systems, MIT.
  • [S101] R. Grosse, “Three sketches of asl-4 safety case components,” Anthropic Alignment Science Blog, 2024.
@@
  • [S102] M. Loba, N. F. Salem, M. Nolte, A. Dotzler, D. Ludwig, and M. Maurer, “Toward a harmonized approach–requirement-based structuring of a safety assurance argumentation for automated vehicles,” arXiv preprint arXiv:2505.03709, 2025.
  • [S103] M. Bagheri, J. Lamp, X. Zhou, L. Feng, and H. Alemzadeh, “Towards developing safety assurance cases for learning-enabled medical cyber-physical systems,” arXiv preprint arXiv:2211.15413, 2022.
  • [S104] M. Balesni, M. Hobbhahn, D. Lindner, A. Meinke, T. Korbak, J. Clymer, B. Shlegeris, J. Scheurer, C. Stix, R. Shah et al., “Towards evaluations-based safety cases for AI scheming,” arXiv preprint arXiv:2411.03336, 2024.
  • [S105] E. Stensrud, T. Skramstad, J. Li, and J. Xie, “Towards goal-based software safety certification based on prescriptive standards,” in 2011 First International Workshop on Software Certification. IEEE, 2011, pp. 13–18.
@@
  • [S106] Z. Chen, Y. Deng, and W. Du, “Trusta: Reasoning about assurance cases with formal methods and large language models,” Science of Computer Programming, vol. 244, p. 103288, 2025.
  • [S107] P. Koopman, “Ul 4600: What to include in an autonomous vehicle safety case,” Computer, vol. 56, no. 05, pp. 101–104, 2023.
  • [S108] Y. Idmessaoud, J.-L. Farges, E. Jenn, V. Mussot, A. F. Pires, F. Chenevier, and R. C. Laguna, “Uncertainty in assurance case pattern for machine learning,” in Embedded Real Time Systems (ERTS), 2024.
@@
  • [S109] J. McDermid, Y. Jia, and I. Habli, “Upstream and downstream AI safety: Both on the same river?” arXiv preprint arXiv:2501.05455, 2024.
  • [S110] K. K. Shahandashti, A. B. Belle, M. M. Mohajer, O. Odu, T. C. Lethbridge, H. Hemmati, and S. Wang, “Using gpt-4 turbo to automatically identify defeaters in assurance cases,” in 2024 IEEE 32nd International Requirements Engineering Conference Workshops (REW). IEEE, 2024, pp. 46–56.
@@
  • [S111] T. Myklebust, T. Stålhane, G. D. Jenssen, and I. Wærø, “Autonomous cars, trust and safety case for the public,” in 2020 Annual Reliability and Maintainability Symposium (RAMS). IEEE, 2020, pp. 1–6.
@@
  • [S112] J. Bragg and I. Habli, “What is acceptably safe for reinforcement learning?” in International Conference on Computer Safety, Reliability, and Security. Springer, 2018, pp. 418–430.